OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–1, 4) and is parallel to the line y = 2x – 3?

OpenStudy (anonymous):

Parallel lines have the same slope. So slope = 2 Enter all the values into the point-slope formula, so y - 4 = 2 ( x -(-1)) y = 2 x +6

OpenStudy (anonymous):

x – 2y = 6 x + 2y = 6 2x – y = –6 2x + y = –6

OpenStudy (anonymous):

Those are my options

OpenStudy (anonymous):

I got that same answer you did but it is not an option

OpenStudy (anonymous):

Just add subtract terms from each side.... y = 2x+6 y-6 = 2x -6 = 2x - y

OpenStudy (anonymous):

sorry....should have read add/subtract terms